eviant:

If f(x)=1/2x-6, what is the equation for f^-1(x)?

eviant:

f(x)=\[\frac{ 1 }{ 2}\]x-6

eviant:

@dude

dude:

\(f(x)=\large \frac{ 1 }{ 2}x-6\) Flip x and y and then solve for y \(x=\large \frac12y-6\)

eviant:

x-6=1/2y

eviant:

@dude

dude:

ADd 6 on both sides

eviant:

x-6+6=1/2y+6 x=1/2y+6

eviant:

@dude

dude:

No no I meant \(x+6=\frac12y\)